# Decoding Visual Experience and Mapping Semantics through Whole-Brain Analysis Using fMRI Foundation Models

### v2

@@ Abstract
Neural decoding, the process of understanding how brain activity corresponds to different stimuli, has been a primary objective in cognitive sciences. Over the past three decades, advancements in functional Magnetic Resonance Imaging (fMRI) and machine learning have greatly improved our ability to map visual stimuli to brain activity, especially in the visual cortex. Concurrently, research has expanded into decoding more complex processes like language and memory across the whole brain, utilizing techniques to handle greater variability and improve signal accuracy. We argue that “seeing” involves more than just mapping visual stimuli onto the visual cortex; it engages the entire brain, as various emotions and cognitive states can emerge from observing different scenes. In this paper, we develop algorithms to enhance our understanding of visual processes by incorporating whole-brain activation maps while individuals are exposed to visual stimuli. We utilize large-scale fMRI encoders and Image generative models (encoders & decoders) pre-trained on large public datasets, which are then fine-tuned through Image-fMRI contrastive learning. Our models hence can decode visual experience across the entire cerebral cortex, surpassing the traditional confines of the visual cortex. Using a public dataset (BOLD5000), we first compare our method with state-of-the-art approaches to decoding visual processing and show improved predictive semantic accuracy by 43%. A network ablation analysis suggests that beyond the visual cortex, the default mode network contributes most to decoding stimuli, in line with the proposed role of this network in sense-making and semantic processing. Additionally, we implemented zero-shot imagination decoding on an extra validation dataset, achieving a p-value of 0.0206 for mapping the reconstructed images and ground-truth text stimuli, which substantiates the model’s capability to capture semantic meanings across various scenarios. These findings underscore the potential of employing comprehensive models in enhancing the nuanced interpretation of semantic processes.
Keywords: Computer Vision, Cognitive Science, Artificial Intelligence, Generative Models
1 Main
@@
Understanding the relationship between brain activity and the characteristics of external stimuli, commonly known as “neural decoding” [26] or identifying “neural representations,” [28] is fundamental to advancing neuroscience. In the last three decades, the development and continuous improvement of functional magnetic resonance imaging (fMRI) allowed researchers to map spatial patterns of activity at the level of millimeter-cubed “voxels” onto patterns of stimuli to understand where in the brain information about different stimuli characteristics is held. The majority of this research has focused on mapping representations of visual stimuli onto the visual cortex with increasing success [26, 53, 10, 49].The field of visual decoding has progressed steadily in terms of the accuracy with which stimuli can be predicted based on brain activity. Still, the pace of this progress has sped up with the recent improvements in machine learning and deep learning methods that can identify more complex relationships between visual activity and visual stimuli [10, 49, 53, 37]. While recent advances in neural decoding have significantly enhanced the ability of models to reconstruct visual stimuli from neural activities in the visual cortex, these models predominantly excel at capturing basic, low-level visual features such as edges, textures, and other pixel-wise attributes [26, 4]. Yet, they struggle to decode the semantics of visual input [10], which is crucial for the understanding of the visual experience.
Running parallel to this work has been attempts to map the representations of more complex higher-order processes such as language [2], semantics [25], movie-watching [20], or autobiographical memory [3] onto either specific networks known to be involved in these processes, or more recently the whole brain [41]. Since these representations vary significantly across individuals both in terms of their functional units and spatial arrangements, these works collect datasets that sample more participants than those used in visual studies and statistical approaches such as multi-voxel pattern analysis [36] or representational similarity analysis [29] are leveraged to capture between-subject representation and variance. Recent evidence increasingly suggests that whole-brain processing plays a crucial role in many fundamental tasks, including visual perception. For example, one study suggests that there is a retinotopic code in the default mode network that allows a mapping between past experiences and incoming visual information [51]. Another study suggests that visual images are encoded in terms of the actions they afford, represented across the whole brain [56]. Predictive coding [1] and active inference [15] theories propose that all visual processing is subject to a comparison with an internal model of the world provided via top-down connections with higher-order brain regions. This suggests that advances in decoding visual activity will benefit from the inclusion of whole-brain information. However, whole-brain data presents a significantly higher dimensionality, requiring orders of magnitude larger datasets to effectively train machine learning algorithms for neural decoding. This is why methods optimized for visual cortex activity have been successful in decoding visual stimuli-the visual cortex comprises a more constrained set of voxels, allowing smaller and less complex models to perform the task effectively. Expanding decoding efforts to whole-brain data necessitates more sophisticated models and larger datasets to capture the complexity of neural processing across multiple brain regions.
@@
Inspired by these two lines of research and making use of recently developed foundation models [5] that leverage pretraining on large datasets to maximize predictive accuracy in more specialized data, we propose a novel approach to decode visual representations present throughout the whole brain (Fig. 1), referred to as Whole-brain Analysis of Visual Experience (WAVE). It leverages the ability of these foundation models [5] to identify between-subjects patterns from large amounts of generic data (e.g., various tasks of fMRI and natural image separately) to improve the predictions of models applied to smaller, more specialized datasets even for tasks that involve large amounts of between-person variance (e.g., mapping semantic representations). First, we compare our method with state-of-the-art approaches to decoding visual processing and show improved predictive accuracy. We highlight that even after removing the visual cortex we can show high predictive accuracy on a traditional visual dataset, emphasizing the importance of whole-brain processing for even simple visual tasks. Subsequently, we enhance our model’s interpretability through post-hoc analyses. Our findings reveal that higher-order cortical hierarchy provides more predictive power for visual scenes with greater complexity. Furthermore, we demonstrate the model’s ability to capture semantic information across different scenarios. We used a separate fMRI imagination dataset where participants gave verbal descriptions followed by fMRI during scenario visualization. Our model excelled at decoding scenarios semantically similar to images from the original dataset. This highlights its capacity to generalize semantic understanding beyond the original training data. This research represents both methodological and conceptual advances: we believe that using these methods can improve the accuracy of brain-stimuli mapping in a way that can be applied across a greater range of datasets. Ultimately, this could lead to more precise modeling of neural processes, gaining novel insights into human cognitive processes.
@@ 2 Results
Figure 1: Model Pipeline and performance comparison with previous work. a, Study Formation: The diagram illustrates our model, WAVE, which reconstructs visual stimuli from fMRI data. b, Model Training Framework: After preprocessing the raw fMRI data, WAVE integrates three modalities: fMRI, image, and text to perform contrastive learning. The features are then passed to a diffusion model for final image reconstruction. c, Brain Saliency Maps: The saliency maps compare the input fMRI data between Mind-Vis and WAVE (ours). Mind-Vis utilizes only visual cortex data (highlighted in red), whereas WAVE employs whole-brain fMRI data. The saliency map for WAVE is based on the model’s attention, demonstrating a broader and more comprehensive engagement of brain regions. d, Quantitative Analysis: This graph compares the performance of the universal model in reconstructing images between Mind-Vis and WAVE. The y-axis represents CLIP accuracy [43] from high-level metrics, and the x-axis represents AlexNet(5) [30] accuracy from low-level metrics [37]. The distributions are displayed on the sides, along with sample images, showcasing the superior performance and robustness of WAVE in capturing both high-level and low-level features.
In this study, we explored human visual experiences through fMRI, leveraging an advanced fMRI foundation model and generative model to reconstruct visual stimuli. Unlike previous studies [49, 10, 52] that predominantly focused on the visual cortex, which is already known for its role in visual processing [24, 23], we extended our method and analysis to encompass the entire brain (Fig. 1c). The generated saliency map highlighted the model’s attention on various brain networks and regions. We employed a range of metrics to evaluate the reconstructed images produced by our model, emphasizing semantic accuracy [10], high-level and low-level correlations [37].
Going beyond the individual model reconstruction typical of previous work [10, 49, 52], we extended our experiment to include a universal setting, wherein a single model decodes neural activities across different subjects neural activities. We conducted qualitative comparison of our approach with another advanced method Mind-Vis in Fig. 1d. We also conducted extensive ablation studies, demonstrating that whole-brain data offers greater semantic advantages over data solely from the visual cortex.
Finally, we analyzed various clusters of images and presented saliency maps from our model, aiding the explainability of our study, and providing insights through cognitive neuroscience statistical analysis.
@@ 2.1 Decoding images from whole-brain fMRIs
We successfully decoded images from whole-brain fMRI data using two distinct methods: building individual models for each subject, as done in visual decoding studies such as [10], and constructing a universal model applicable across four subjects (details in Section 4.1). We evaluated our model’s performance by comparing the reconstructed images against the original visual stimuli, the ground-truth, using a range of metrics that examined the semantic meaning, high-level relationships, and low-level pixel-wise distances (details in Section 4.2).
@@
As shown in Table 1, we benchmarked our model, WAVE, with best in class models MindEye [49] and MindVis [10]. The results of individual subject-based models are detailed in Appendix A.1. In both individual and universal settings, WAVE achieved the best performance in Semantic Top-1 Accuracy and high-level metrics. In low-level metrics, WAVE also outperformed in AlexNet(2) and AlexNet(5). Notably, while SSIM and PixCorr metrics measure pixel-wise correlations, all other metrics from both low-level and high-level categories measure encoded image features from models (e.g. AlexNet, CLIP). (See metrics details in Section 4.2) This suggests that our model excels in semantic and high-level image reconstruction. We also provide rich reconstructed image examples comparing WAVE with MindEye and Mind-Vis in Fig. 2a.
@@
Furthermore, the results from the universal decoding setting, as shown in Table 2 underscore WAVE’s superior performance and the advantage of utilizing whole-brain data compared to Mind-Vis [10]. Due to individual variability, brain fMRI data differ significantly between voxels. The same registered voxel in standard space usually does not have the same function across subjects. However, brain regions tend to maintain similar functions across individuals. Our brain-wide network preprocessed fMRI data showed more advantages in universal settings. Additionally, by leveraging the pre-trained foundation fMRI model, WAVE demonstrated better generalization power, highlighting the robustness and adaptability of our approach. The reconstructed examples in Fig. 2b clearly show our model’s consistency of the generated image across four subjects. WAVE can capture the semantic meaning from brain-wide network activity accurately compare with other methods.
@@
Table 1: An evaluation of our model’s capability to reconstruct images from fMRI data. Employing Semantic metrics as Chen et al. in [10], we adopt a 50-way top-1 accuracy approach (chance=2%) over 1000 trials to gauge the fidelity of reconstructed images against the original stimuli. We used a similar setting to previous study, using Low-Level and High-Level metrics. We used pixel-wise correlation (PixCorr) and structural similarity index metrics (SSIM). Additionally, we assess correlation distance using the EfficientNet-B1 (Eff) [54] and SwAV-ResNet50 (SwAV) [8] models. For other metrics, two-way identification (chance=50%) is used for evaluation, consistent with methodologies described by Ozcelik and VanRullen, Scotti et al.. Two experimental setups were investigated: one involving individualized models for each subject, with an average taken across four subjects’ metric results (detailed results in the appendix), and another employing a universal model for all subjects.
@@
Setting Model Low-Level High-Level Semantic PixCorr↑ SSIM↑ AlexNet(2)↑ AlexNet(5)↑ Incep↑ CLIP↑ Eff↓ SwAV↓ Top-1 Acc↑ Individual WAVE (Ours) 0.062 0.199 71.61% 79.70% 68.76% 81.54% 0.892 0.552 25.24% MindEye [49] 0.056 0.217 71.11% 77.31% 66.32 69.98% 0.916 0.592 16.85% Mind-Vis [10] 0.074 0.308 67.42% 72.74% 66.97% 70.07% 0.914 0.556 17.66% Universal WAVE (Ours) 0.050 0.194 69.47% 78.31% 68.41% 78.41% 0.902 0.591 20.75% Mind-Vis [10] 0.036 0.272 60.56% 68.92% 64.25% 64.47% 0.938 0.593 9.19%
@@
As shown in Table 2, whole-brain network data demonstrated significantly better semantic characteristics compared to data derived from visual cortex voxels. In both individual and universal settings, whole-brain fMRI data achieved state-of-the-art results across all high-level metrics. While PixCorr and SSIM metrics relate more to visual cortex area features, the whole-brain data incorporates complementary information from high-hierarchy functional brain regions involved in visual processing. This comprehensive approach highlights the semantic advantages of using whole-brain data for image reconstruction.
@@
Setting Data Low-Level High-Level Semantic PixCorr↑ SSIM↑ AlexNet(2)↑ AlexNet(5)↑ Incep↑ CLIP↑ Eff↓ SwAV↓ Accuracy↑ Indvidual Visual Cortex 0.056 0.217 71.11% 77.31% 66.32% 68.11% 0.916 0.592 16.85% Whole Brain 0.084 0.215 71.48% 79.49% 68.64% 76.29% 0.908 0.579 19.60% Universal Visual Cortex 0.081 0.214 68.32% 73.82% 63.31% 69.75% 0.929 0.609 14.43% Whole Brain 0.091 0.213 68.96% 76.67% 66.60% 76.18% 0.903 0.596 18.72%
@@
Figure 2: Comparison of fMRI data decoding using WAVE and other advanced methods. a, Reconstructed images using individual model settings for WAVE (Whole Brain), MindEye, and Mind-Vis across the fMRI visual cortex. b, Universal settings across all four subjects in the BOLD5000 dataset. This panel demonstrates the generalization capability of WAVE, MindEye, and Mind-Vis among different subjects. c, The saliency map of the WAVE model using data where the visual cortex has been masked, showing the top 20 regions of interests. d, Decoding from the visual cortex-masked model: examples of reconstructed images using non-visual fMRI regions. In each pair, the left image shows the original visual stimulus, and the right image shows the reconstructed image generated by our visual cortex-masked method (WAVE). e, Impact of network ablation on Decoding Accuracy (subject CSI-1). The left side displays generated images resulting from the masking of each of the seven networks. The right side features a box plot illustrating the decoding accuracy for each network ablation, highlighting the accuracy reduction when specific networks are removed. The red dashed line above represents the whole-brain decoding accuracy for comparison.
The guiding principle of our research is that brain regions beyond the visual cortex contain valuable information crucial for decoding visual experiences. We posit that these non-visual regions primarily encode the semantic aspects of visual stimuli, thus providing essential information for semantic decoding tasks and image reconstruction. To substantiate this claim, we conducted several experiments outlined below:
  • Semantic decoding without visual network. We utilized the WAVE model to reconstruct images without data from the visual network, as shown in Figure 2d. These reconstructions retained high quality, with the average semantic accuracy across four subjects reaching 18.98%. In contrast, whole-brain decoding achieved a higher average accuracy of 25.25%, as presented in Table 1. The high quality reconstruction beyond visual cortex suggests non-visual network also includes significant information for visual experiences decoding. Additionally, the saliency map of the top 20 regions for non-visual regions decoding, is displayed in Figure 2c.
  • Network mask impact ablation: To evaluate the contribution of the visual network to decoding, we conducted a network ablation study, systematically masking one of the Yeo 7 networks [63] at a time and using the remaining six for decoding. The results for subject CSI-1, depicted in Figure 2e, show that the removal of the visual network results in a significant accuracy reduction, marked by a red dashed line, compared with whole-brain decoding. In contrast, masking any of the other six networks did not affect semantic accuracy to the same extent as masking the visual network. However, the results also revealed that the default mode and dorsal attention networks play important roles in visual decoding, as their removal also resulted in considerable accuracy reductions. Further reconstructed examples of individual network masking are also provided in Figure 2e. This is in line with expectations, given that the dorsal attention network is adjacent to the visual network and performs higher order visual functions, and the default mode network has a key role in semantic processing and “making sense” of visual input [64].

### v3

@@ Abstract
Neural decoding, the process of understanding how brain activity corresponds to different stimuli, has been a primary objective in cognitive sciences. Over the past three decades, advances in functional Magnetic Resonance Imaging (fMRI) and machine learning have greatly improved our ability to map visual stimuli to brain activity, especially in the visual cortex. Concurrently, research has expanded to decode more complex processes, such as language and memory across the whole brain, using techniques to handle greater variability and improve signal accuracy. We argue that “seeing” involves more than just mapping visual stimuli onto the visual cortex; it engages the entire brain, as various emotions and cognitive states can emerge from observing different scenes. In this paper, we develop algorithms to enhance our understanding of visual processes by incorporating whole-brain activation maps while individuals are exposed to visual stimuli. We utilize transformer-based large-scale fMRI encoders and Image generative models (encoders & decoders) pre-trained on large public datasets, which are then fine-tuned through Image-fMRI contrastive learning. Our models can decode visual experience across the entire cerebral cortex, surpassing the traditional confines of the visual cortex. Using a public dataset (BOLD5000), we first compare our method with state-of-the-art approaches for decoding visual processing and show improved predictive semantic accuracy by 43%. A network ablation analysis suggests that, beyond the visual cortex, the default mode network contributes significantly to stimulus decoding, in line with the proposed role of this network in sense-making and semantic processing. Additionally, we implemented zero-shot (i.e., no additional training) imagination decoding on an independent validation dataset, achieving a statistically significant ($p$-value <0.05<0.05) mapping between the reconstructed images and ground-truth text stimuli. This demonstrates the model’s generalization capability to capture semantic meanings across various scenarios. These findings underscore the potential of employing comprehensive models to enhance the nuanced interpretation of semantic processes.
keywords:
Neural Decoding , Foundation Models , Functional MRI , Visual Reconstruction
journal: Medical Image Analysis
[…]
[aff1]organization=Department of Psychiatry and Behavioral Sciences, Stanford University, city=Stanford, state=CA, country=USA
[…]
[aff2]organization=Department of Computer Science, Stanford University, city=Stanford, state=CA, country=USA
[…]
[aff3]organization=Department of Neurobiology, University of Chicago, city=Chicago, state=IL, country=USA
[…]
[aff4]organization=UST, city=Tampa, state=FL, country=USA
[…]
[aff5]organization=Google DeepMind, city=Toronto, state=Ontario, country=Canada
1 Introduction
@@
Understanding the relationship between brain activity and the characteristics of external stimuli, commonly known as “neural decoding” [Kamitani and Tong, 2005] or identifying “neural representations,” [Kriegeskorte and Diedrichsen, 2019] is fundamental to advancing neuroscience. In the last three decades, the development and continuous improvement of functional magnetic resonance imaging (fMRI) has allowed researchers to map spatial patterns of activity at the level of millimeter-cubed “voxels” onto patterns of stimuli to discern how specific information is encoded. The majority of research has focused on mapping representations of visual stimuli onto the visual cortex with increasing success Kamitani and Tong [2005], Takagi and Nishimoto [2023], Chen et al. [2023], Scotti et al. [2024], Xia et al. [2024].
The field of visual decoding has advanced steadily in terms of the accuracy with which stimuli can be predicted from brain activity. This progress has accelerated with the recent improvements in machine learning and deep learning methods that can identify more complex non-linear relationships between visual activity and visual stimuli [Chen et al., 2023, Scotti et al., 2024, Takagi and Nishimoto, 2023, Ozcelik and VanRullen, 2023].
While recent advances in neural decoding have significantly enhanced the ability to reconstruct visual stimuli from the visual cortex, these models predominantly excel at capturing basic, low-level visual features such as edges, textures, and other pixel-wise attributes [Kamitani and Tong, 2005, Beliy et al., 2019]. However, human vision is a holistic process where sensory input is integrated with memory and emotion Pessoa and Adolphs [2010], Khan et al. [2011].
Running parallel to this work have been attempts to map the representations of more complex higher-order processes such as language [Anderson and Lin, 2019], semantics [Huth et al., 2016], movie-watching [Guntupalli et al., 2016], or autobiographical memory [Anderson et al., 2020]. Unlike localized visual decoding, these studies demonstrate that information is often encoded across broad cortical networks Power et al. [2017]. For instance, seminal work by Huth et al. [Huth et al., 2016] revealed that semantic concepts are mapped systematically across the cerebral cortex, suggesting that “meaning” is not confined to a single region. However, because these distributed patterns exhibit high inter-subject variability, traditional statistical approaches, such as multi-voxel pattern analysis (MVPA) [Norman et al., 2006] or representational similarity analysis (RSA) [Kriegeskorte et al., 2008], often require datasets with a large amount of participants, taking advantage of individual differences to understand shared variance. This can be contrasted with visual decoding that typically uses longer scans in fewer individuals to capture basic visual processes that vary less across people.
Recent evidence increasingly implicates the whole-brain in fundamental visual tasks. Notably, studies suggest that the default mode network (DMN) may maintain a reinotopic code to bridge past experiences with incoming visual information, while affordance-based representations of objects appear distributed globally  [Steel et al., 2024, Thornton and Tamir, 2024]. Theories of predictive coding [Aitchison and Lengyel, 2017] and active inference  [Friston et al., 2009] further argue that visual perception relies on top-down comparisons with internal world models housed in higher-order regions. Despite this theoretical motivation, expanding visual decoding to the whole-brain presents a severe computational bottleneck due to the type of data typically used for these tasks.” Whole-brain data introduces a feature space that vastly outscales the number of available training samples and the lack of between-subject variance in less basic visual processes (e.g., those involving subjective experience). Previous methods circumvented this by restricting analyses to the visual cortex, allowing simpler linear models to function effectively within a constrained search space Takagi and Nishimoto [2023], Kamitani and Tong [2005] but only allowing the modeling of basic processes that are similar across participants. To decode more complex, variable visual processes from the whole brain without overfitting, we therefore require models with strong inductive priors and the capacity to transfer robust between-subject representations to high-dimensional, noisy inputs.
@@
To address this challenge, we introduce WAVE (Whole-brain Analysis of Visual Experience), a framework that leverages the capacity of foundation models [Bommasani et al., 2021] to extract universal neural dynamics from large-scale generic data. By transferring these learned priors to the visual decoding domain, WAVE can effectively model high-dimensional whole-brain representations even in datasets with a limited number of participants (Fig. 1). Our contribution is threefold. First, we compare our method with state-of-the-art approaches to decoding visual processing and show improved predictive accuracy. We highlight that even after removing the visual cortex we can show high predictive accuracy on a traditional visual dataset, emphasizing the importance of whole-brain processing for even simple visual tasks. Subsequently, we enhance our model’s interpretability through post-hoc analyses. Our findings reveal that the higher-order cortical hierarchy provides more predictive power for visual scenes with greater complexity. Furthermore, we demonstrate the model’s ability to capture semantic information across different scenarios. We used a separate fMRI imagination dataset where participants gave verbal descriptions followed by fMRI during scenario visualization. Our model excelled at decoding scenarios semantically similar to images from the original dataset. This highlights WAVE’s capacity to generalize semantic understanding beyond the original training data. This research represents both methodological and conceptual advances: we believe that our method can improve the accuracy of brain-stimuli mapping in a way that can be applied across a greater range of datasets. Ultimately, this could lead to more precise modeling of distributed neural processes, gaining novel insights into human cognition.
@@ 2 Results
Figure 1: Model Pipeline and performance comparison with previous work. a, Study Formation: The diagram illustrates our model, WAVE, which reconstructs visual stimuli from fMRI data. b, Model Training Framework: After preprocessing the raw fMRI data, WAVE integrates three modalities: fMRI, image, and text to perform contrastive learning. The features are then passed to a diffusion model for final image reconstruction. c, Brain Saliency Maps: The saliency maps compare the input fMRI data between Mind-Vis and WAVE. Mind-Vis utilizes only visual cortex data (highlighted voxels in red), whereas WAVE employs whole-brain fMRI data. The saliency map for WAVE is based on the model’s attention, demonstrating a broader and more comprehensive engagement of brain regions. d, Quantitative Analysis: This graph compares the performance of the universal model in reconstructing images between Mind-Vis and WAVE. The y-axis represents CLIP accuracy [Radford et al., 2021] from high-level metrics, and the x-axis represents AlexNet(5) [Krizhevsky et al., 2012] accuracy from low-level metrics [Ozcelik and VanRullen, 2023]. The distributions are displayed on the sides, along with sample images, showing the superior performance of WAVE in capturing both high-level and low-level features.
In this study, we explored human visual experiences through fMRI, leveraging an architecture that integrates a transformer-based fMRI foundation model with a latent diffusion generative model. While prior studies Scotti et al. [2024], Chen et al. [2023], Sun et al. [2024], Xia et al. [2024] predominantly focused on the visual cortex, which is already known for its role in visual processing Huff et al. [2018], Hubel and Wiesel [1964], this whole-brain approach allowed us to generate attention-based saliency maps that reveal how semantic information is encoded across diverse brain networks beyond early visual areas (Fig. 1c). We employed a range of metrics to evaluate the reconstructed images produced by our model, emphasizing semantic accuracy Chen et al. [2023], high-level and low-level correlations Ozcelik and VanRullen [2023].
Going beyond the individual model reconstruction typical of previous work Chen et al. [2023], Scotti et al. [2024], Sun et al. [2024], we extended our experiment to include a universal setting, wherein a single model decodes neural activities across different subjects. We conducted a qualitative comparison of our approach with another advanced method Mind-Vis in Fig. 1d. We also conducted extensive ablation studies, demonstrating that whole-brain data offers greater semantic advantages over data solely from the visual cortex.
Finally, we analyzed various clusters of images and presented brain saliency maps from our model, improving the explainability of our method and providing insights through cognitive neuroscience statistical analysis.
@@ 2.1 Decoding images from whole-brain fMRIs
We successfully decoded images from whole-brain fMRI data using two distinct methods: building individual models for each subject, as done in visual decoding studies such as Chen et al. [2023], and constructing a universal model applicable across four subjects (details in Section 4.1). We evaluated our model’s performance by comparing the reconstructed images against the original visual stimuli, the ground-truth, using a range of metrics that examined the semantic meaning, high-level relationships, and low-level pixel-wise distances (details in Section 4.2).
@@
As shown in Table 1, we benchmarked our model, WAVE, with best in class models MindEye Scotti et al. [2024] and MindVis Chen et al. [2023]. The results of individual subject-based models are detailed in Table 4. In both individual and universal settings, WAVE achieved the best performance in Semantic Top-1 Accuracy and high-level metrics. In low-level metrics, WAVE also outperformed AlexNet(2) and AlexNet(5). Notably, while SSIM and PixCorr metrics measure pixel-wise correlations, all other metrics from both low-level and high-level categories measure encoded image features from models (e.g. AlexNet, CLIP). (See metrics details in Section 4.2) This suggests that our model excels in semantic and high-level image reconstruction. We also provide rich reconstructed image examples comparing WAVE with MindEye and Mind-Vis in Fig. 2a.
@@
Furthermore, the results from the universal decoding setting, as shown in Table 2 underscore WAVE’s superior performance and the advantage of utilizing whole-brain data compared to Mind-Vis Chen et al. [2023]. Due to individual variability, brain fMRI data differ significantly between voxels. The same registered voxel in standard space usually does not have the same function across subjects. However, brain regions tend to maintain similar functions across individuals. Our brain-wide network preprocessed fMRI data showed more advantages in universal settings. Additionally, by leveraging the pre-trained foundation fMRI model, WAVE demonstrated better generalization power, highlighting the robustness and adaptability of our approach. The reconstructed examples in Fig. 2b clearly show our model’s consistency of the generated image across four subjects. WAVE can capture the semantic meaning from brain-wide network activity accurately compared with other methods.
@@
Table 1: An evaluation of our model’s capability to reconstruct images from fMRI data. Employing Semantic metrics as  Chen et al. [2023], we adopt a 50-way top-1 accuracy approach (chance=2%) over 1000 trials to gauge the fidelity of reconstructed images against the original stimuli. We used a similar setting to previous study, using Low-Level and High-Level metrics. We used pixel-wise correlation (PixCorr) and structural similarity index metrics (SSIM). Additionally, we assess correlation distance using the EfficientNet-B1 (Eff) [Tan and Le, 2019] and SwAV-ResNet50 (SwAV) [Caron et al., 2020] models. For other metrics, two-way identification (chance=50%) is used for evaluation, consistent with methodologies described by Ozcelik and VanRullen [2023], Scotti et al. [2024]. Two experimental setups were investigated: one involving individualized models for each subject, with an average taken across four subjects’ metric results (detailed results in the appendix), and another employing a universal model for all subjects.
@@
Setting Model Low-Level High-Level Semantic PixCorr↑ SSIM↑ AlexNet(2)↑ AlexNet(5)↑ Incep↑ CLIP↑ Eff↓ SwAV↓ Top-1 Acc↑ Individual WAVE (Ours) 0.062 0.199 71.61% 79.70% 68.76% 81.54% 0.892 0.552 25.24% MindEye [Scotti et al., 2024] 0.056 0.217 71.11% 77.31% 66.32 69.98% 0.916 0.592 16.85% Mind-Vis [Chen et al., 2023] 0.074 0.308 67.42% 72.74% 66.97% 70.07% 0.914 0.556 17.66% Universal WAVE (Ours) 0.050 0.194 69.47% 78.31% 68.41% 78.41% 0.902 0.591 20.75% Mind-Vis [Chen et al., 2023] 0.036 0.272 60.56% 68.92% 64.25% 64.47% 0.938 0.593 9.19%
@@
As shown in Table 2, whole-brain network data demonstrated significantly better semantic characteristics compared to data derived from visual cortex voxels. In both individual and universal settings, whole-brain fMRI data achieved state-of-the-art results across all high-level metrics. While PixCorr and SSIM metrics relate more to visual cortex area features, the whole-brain data incorporates complementary information from high-hierarchy functional brain regions involved in visual processing. This comprehensive approach highlights the semantic advantages of using whole-brain data for image reconstruction. Furthermore, additional analyses on the encoder’s representational capacity (Table 8 and Table 9) suggest that the pre-trained fMRI model plays a pivotal role in effectively extracting and integrating these complex whole-brain signals.
@@
[…]
Figure 2: Comparative Evaluation and Network Ablation of the WAVE Framework. a, Qualitative comparison of reconstructed images. Columns display the original visual stimulus alongside reconstructions from WAVE and baselines. b, Universal settings across all four subjects in the BOLD5000 dataset. This panel demonstrates the generalization capability of WAVE, MindEye, and Mind-Vis among different subjects. c, The saliency map of the WAVE model using data where the visual cortex has been masked, showing the top 20 regions of interests. d, Decoding from the visual cortex-masked model: examples of reconstructed images using non-visual fMRI regions. In each pair, the left image shows the original visual stimulus, and the right image shows the reconstructed image generated by our visual cortex-masked method (WAVE). e, Impact of network ablation on Decoding Accuracy (subject CSI-1). The left side displays generated images resulting from the masking of each of the seven networks. The right side features a box plot illustrating the decoding accuracy for each network ablation, highlighting the accuracy reduction when specific networks are removed. The red dashed line above represents the whole-brain decoding accuracy for comparison.
@@
Our research is guided by the hypothesis that neural representations of visual experience are distributed well beyond the primary visual cortex. We posit that non-visual regions encode essential high-level semantic aspects of visual stimuli, which are crutial for accurate semantic-correct image reconstruction. To validate this, we performed a series of computational lesioning experiments to isolate the contributions of specific brain networks.
First, we assessed the model’s capacity to decode visual stimuli solely from non-visual regions by completely masking the visual network. As illustrated in Figure 2d, the resulting reconstructions retained remarkable semantic fidelity, achieving an average semantic accuracy of 18.98% across four subjects. While this is lower than the 25.24% accuracy achieved by the whole-brain model (Table 1), the relatively high performance confirms that non-visual networks contain sufficient information to drive semantic decoding. The saliency map highlighting the top 20 non-visual regions contributing to this process is displayed in Figure 2c.
To further dissect these contributions, we conducted a systematic network ablation study using the Yeo 7-network parcellation [Yeo et al., 2011]. We iteratively masked one network at a time while using the remaining six for decoding. Results for subject CSI-1 (Figure 2e) indicate that while removing the visual network caused the most significant drop in accuracy (marked by the red dashed line), the exclusion of the Default Mode Network (DMN) and Dorsal Attention Network also resulted in substantial performance decrements. This suggests these networks are not merely auxiliary but integral to the decoding process. This finding aligns with established neuroscientific literature: the dorsal attention network supports higher-order visual attention, while the DMN is implicated in integrating semantic context and “making sense” of sensory inputs Yeshurun et al. [2021].